Two objects A and B having inertias Ma=m and Mb=1.1m are moving with velocities Va=V and Vb=8.4v.
Find the velocities of object B in the zero momentum frame of reference
1
Expert's answer
2020-05-19T10:53:06-0400

The velocity of center of mass

"V=\frac{M_av_a+M_bv_b}{M_a+M_b}\\
=\frac{mv+1.1m\times 8.4v}{m+1.1m}=4.9v"

Hence, the velocities of object B in the zero momentum frame of reference


"v_b'=v_b-V=8.4v-4.9v=3.5v"